Please help me solve this problem: (Mixture Problem)

1. How many grams of Brand A coffee worth P4.00 per gram must be added to 100 grams of Brand B coffee worth P7.00 per gram to produce a mixture worth P6.00 per gram?

Answer by RAY100(1637)

A g (4 #/g) + 100 gms (7 #/g) = (A + 100)g * 6 #/g
.
4A +700 = 6 (A+100)= 6A + 600
.
100 = 2A
.
50 g = A,,,,,,answer
.
to check,,50*4 +100*7 = (100+50)*6
.
200 +700 =900 = 150 * 6 ,,,,ok
.